determine the work done by a force F = xyi + yzj + xzk in taking a particle along the path defined by the equation r(t)=ti + 2t²j + t³k, 0≤t≤1 from t=0 to t=2.


1
Expert's answer
2021-06-02T13:07:38-0400

Solution:-

we have the following data given

Force "F = xy\hat{i} + yz\hat{j} + xz\hat{k }"

Path equation "r(t) = t\hat{i} + 2t^2\hat{j} + t^3\hat{k}"

"d(r) = \hat{i}+4t\hat{j}+3t^2\hat{k}dt"


Work done "W=\int F.dr"

"W=\int(xy\hat{i} + yz\hat{j} + xz\hat{k })(\hat{i} +4t\hat{j} + 3t^2\hat{k})dt"

"=\int xydt+yz\times 4t dt+xz\times 3t^2 dt"

Now, x = t, "y=t^2" and "z=t^3"

"=\int t^3dt + 4t^6 dt+3t^6 dt"

"=\int 7t^6 dt + \int t^3 dt"

"=t^7+\frac{t^4}{4}"

Now, substituting the values of t = 1

"=\frac{5}{4} \ answer"